Box A, resting on a ramp, is connected box B through a pulley system. The coefficient of kinetic friction for box A, the height of box B above the ground, and the angle of inclination for the ramp are provided in the information section below. After the boxes are released, how far up the ramp will Box A travel?


m A = 7.0 kg

m B = 4.0 kg

µ kA = 0.10

h = 1.5 m

θ = 20°


1
Expert's answer
2021-04-06T18:25:12-0400

"m_Agsin\theta-T+F_f=m_Aa,"

"N-m_Agcos\theta=0,"

"T-m_Bg=m_Ba,"

"F_f=\mu N,"

"N=m_Agcos \theta,"

"T=m_B(g+a),"

"a=\frac{g(m_Asin \theta-m_B+\mu m_Acos 
 \theta }{m_A+m_B}=-0.85~\frac{m}{s^2}<0,"

"l=\frac{h}{sin\theta}=4.39~m" down.